OpenStudy (anonymous):

find the term independent of x in the expression of (6x+1/x)^12

OpenStudy (anonymous):

12C6 (6x)^6 (1/x)^6 = 12C6 6^6

OpenStudy (anonymous):

- which is a huge number!

OpenStudy (anonymous):

can u plz explain little bit as i m non maths student

OpenStudy (anonymous):

the formula used is (r + 1) term of (x + b)^n = nCr x^(n-r) b^r

OpenStudy (anonymous):

nCr means the number of combinations of r in n for example, 5C2 = 5*4 --- = 10 2*1

OpenStudy (anonymous):

the expansion is called the binomial theorem a simple example is (x+3)^3 = x^3 + 3x^2*3 + 3x*3^2 + 9

OpenStudy (anonymous):

typo - last term is 3^3 = 27